OpenStudy (anonymous):

Andrea can ride her bike 2 mi/h faster than ruby. It takes Andrea 48 min less the travel 50 mi than it does ruby. What is Andrea's rate in mi/h?

OpenStudy (anonymous):

-__-... she doesn't even show the steps. I need steps

OpenStudy (ajprincess):

Let us assume that the speed of Andrea to be x. Then the speed of Ruby will be x-2. \(\large speed=\LARGE\frac{distance}{time}\) Say the time taken by Andrea to cover 50 miles is t. \(\large x=\LARGE\frac{50}{t}\) \(50=xt\) The time taken by Ruby to cover the same 50 miles=t+48 \(\large x-2=\LARGE\frac{50}{t+48}\) \(50=(x-2)(t+48)\) \(xt=(x-2)(t+48)\) Can u find x nw?

OpenStudy (ajprincess):

@steeben can u do?